How can you factor this expression:
(1-square root of 2+square root of 5)(1+square root of 2-square root of 5)

Answer by stanbon(75887)

(1-square root of 2+square root of 5)(1+square root of 2-square root of 5)

(1 - (sqrt2 - sqrt5))(1 + (sqrt2 - sqrt5))
This is the form (a-b)(a+b) which gives the form a^2-b^2
= 1-(sqrt2-sqrt5)^2
=1-[2-2sqrt10+5]
=1-7+2sqrt10
=-6+2sqrt10
Cheers,
Stan H.